A firm has EBIT of $500,000, interest expenses of $300,000, and a cooperate tax rate of 35%.

  1. What is its net income?
  2. What would its net income be if it did not have any debt (and, consequently, no interest expense)?
  3. What are the firm’s interest tax savings?

Indicate the detailed steps on how to use a FINANCIAL CALCULATOR to solve the problems.
Answer #1

EBIT = $500,000

Interest Expense = $300,000

Tax Rate = 35%

1.

Net Income = (500,000 - 300,000)(1 - 0.35)

Net Income = $130,000

2.

If Debt = 0,

Interest Expense = 0

Net Income = (1 - 0.35)(500,000)

Net Income = $325,000

3.

Interest Tax Saving = 0.35(Interest Expense)

Interest Tax Saving = 0.35(300,000)

Interest Tax Saving = $105,000